# MCTED: A Machine-Learning-Ready Dataset for Digital Elevation Model Generation From Mars Imagery

### v1

@@ 1 Introduction
  • We make the source code for dataset cleaning, generation, model training and evaluation open source, pending an ESA licensing process, under the following address: https://github.com/ESA/MCTED.
@@ Acknowledgments
We would like to acknowledge and express our gratitude to the creators of the Day et al. repository for generating high-quality data and making it openly available. This open-access nature of the data repository is invaluable for the scientific community to further analyse, process and develop new methodologies. Such contributions greatly enhance the accessibility of research; therefore, we are making the MCTED dataset, together with all code necessary for its generation, fully open-source, pending an ESA licensing process.
@@ A.3 Verticalisation
To compute the image rotation needed to verticalise the image, two reference points are computed: (i) the first non-black pixel encountered while scanning down the leftmost column, giving its row coordinate $y_{\ell}$, and (ii) the first non-black pixel encountered while scanning rightwards along the bottom row, giving its column coordinate $x_{b}$. Together with the fixed bottom-left corner (0,H1)(0,H-1) these points define a right-angled triangle whose interior angle at (0,H1)(0,H-1) is
@@
The image is clockwise if $y_{\ell}<H/2$ and counter-clockwise otherwise; a horizontal mirror is applied in the former case so that rotation is always carried out in a consistent sense. The image, the DEM and all auxiliary masks are then rotated by $\alpha$ using a rigid Euclidean transform from scikit-image and cv2 for any resampling.
After rotation, a second sweep to eliminate black pixel rows/columns is executed with $\tau=0.1$, so a line is cropped when at least ten per cent of its pixels are black. This is usually aggressive enough to eliminate the thin wedges of near-black padding introduced by interpolation, yet conservative enough to preserve genuine terrain in low-contrast scenes.
The end product is a verticalised, border-free image and DEM.

### v2

@@ 1 Introduction
  • We make the source code for dataset cleaning, generation, model training and evaluation open source under the following address: https://github.com/ESA/MCTED.
@@ Acknowledgments
We would like to acknowledge and express our gratitude to the creators of the Day et al. repository for generating high-quality data and making it openly available. This open-access nature of the data repository is invaluable for the scientific community to further analyse, process and develop new methodologies. Such contributions greatly enhance the accessibility of research; therefore, we are making the MCTED dataset, together with all code necessary for its generation, fully open-source.
@@ A.3 Verticalisation
To compute the image rotation needed to verticalise the image, two reference points are computed: (i) the first non‑black pixel encountered while scanning down the leftmost column, giving its row coordinate $y_{\ell}$, and (ii) the first non‑black pixel encountered while scanning rightwards along the bottom row, giving its column coordinate $x_{b}$. Together with the fixed bottom‑left corner (0,H1)(0,H-1) these points define a right‑angled triangle whose interior angle at (0,H1)(0,H-1) is
@@
The image is clockwise if $y_{\ell}<H/2$ and counter‑clockwise otherwise; a horizontal mirror is applied in the former case so that rotation is always carried out in a consistent sense. The image, the DEM and all auxiliary masks are then rotated by $\alpha$ using a rigid Euclidean transform from scikit-image and cv2 for any resampling.
After rotation, a second sweep to eliminate black pixel rows/columns is executed with $\tau=0.1$, so a line is cropped when at least ten per cent of its pixels are black. This is usually aggressive enough to eliminate the thin wedges of near‑black padding introduced by interpolation, yet conservative enough to preserve genuine terrain in low‑contrast scenes.
The end product is a verticalised, border‑free image and DEM.